### v2

# WiFo-M2: Empower Wireless Communications With Plug-and-Play Environment Sensing via Foundation Model

Abstract
The emerging convergence of next-generation wireless networks and agentic artificial intelligence (AI) is inspiring a new vision: embodied intelligent network entities utilize environmental sensing to refine their physical-layer (PHY) actions. Despite a growing body of preliminary work, prevailing small and task-specific AI models require extensive manual design of data pre-processing, network architecture, and fine-tuning, leaving them tightly coupled to particular PHY actions, system configurations, and deployment scenarios. To address this, we propose a paradigm shift with WiFo-M2, a foundation model that enables environment sensing to be easily integrated into PHY actions, delivering universal performance gains. To extract generalizable out-of-band (OOB) channel-aware features from environment sensing, we introduce ContraSoM, a contrastive pre-training strategy. Once pre-trained, WiFo-M2 infers future OOB channel-aware features from historical sensory data and strengthens feature robustness via modality-specific data augmentation. Experiments show that WiFo-M2 improves the performance of diverse PHY actions and demonstrates strong generalization to unseen scenarios.
I Introduction
The 2030 era of networking will feature hundreds of billions of connected devices, immersive applications, and tight cyber physical loops that demand extreme reliability, ultra low latency, and sustainable efficiency [1]. Meeting these requirements calls for Agentic AI as a foundational paradigm in which autonomous networked agents perceive, reason, plan, and act [2]. Today, the application of Agentic AI in wireless networks is rapidly expanding, providing a robust framework for autonomous optimization and self-evolving in increasingly complex and dynamic environments [2]. Against this backdrop, Cheng et al. recently introduced Embodied Intelligent Wireless (EIW) [3] and envisioned it as the new paradigm for next-generation wireless communications. Under EIW, conventional network nodes such as base stations (BSs) and user equipment (UEs) evolve into intelligent network entities endowed with observation–decision–action capabilities.
Driven by the rapid advancement of embodied intelligence, intelligent embodied network entities are increasingly equipped with comprehensive environmental sensing capabilities [4]. Motivated by this trend, the wireless communications community is actively exploring how these ubiquitous sensing sources can be leveraged to enhance physical layer (PHY) actions. Here, “actions” concretely refers to transceiver modules, e.g., channel estimation/prediction, and beamforming. To systematically give the design principle and underlying mechanisms of such sensing-refined PHY actions, Cheng et al. proposed the first framework of this field, Synesthesia of Machines (SoM) [4]. To date, numerous studies have investigated promising applications, such as using multi-modal sensing to boost channel estimation acquisition performance [7, 5, 6], to streamline beam selection and training [8, 9, 10, 11], and to optimize precoding [12].
However, as illustrated in Fig. 1, the field of sensing‑refined PHY actions remains dominated by traditional per‑task “small models” that deliver isolated gains but generalize poorly. They are scenario‑bound and configuration‑locked, lacking a cohesive mechanism to convert rich observations into consistently better PHY actions across environments. Moreover, the per-task paradigm keeps the research scope narrow and non-transferable. Each PHY action must be served by a customized data pipeline and model, driving up engineering cost and stalling large-scale deployment. Consequently, existing studies mostly address a few vision-friendly tasks such as beam prediction and handover, whereas broader needs like channel interpolation or prediction receive little attention. Finally, current BSs, though equipped with multi-modal sensors, still lack agentic‐AI capabilities. They passively fit sensing inputs to predefined outputs rather than learning latent scene dynamics and generally useful channel-aware features, leaving them unable to exploit observations under occlusion or to anticipate fine-grained out-of-band (OOB) features.
Figure 1: Comparisons between conventional RF-based/sensing-refined wireless communications paradigm, and the proposed FM-based approach of refining PHY actions with environment sensing under EIW paradigm.
To propel current sensing-refined PHY research toward the EIW paradigm and the vision of Agentic AI, we advocate a foundation model (FM) based approach that equips intelligent entities with a universal, task-agnostic observation-to-action bridge. In this paper, we introduce WiFo-M2, the first universal FM that transforms multi-modal sensing into OOB channel-aware features, thereby serving as a plug-and-play performance booster across diverse PHY actions, as shown in Fig. 1. To establish one-to-one positive sample pairs between channel state information (CSI) and multi-modal sensory data, we first develop a unified multi-modal sensing processing pipeline. Then, we design the ContraSoM pre-training strategy, which is designed to align the high-dimensional representations extracted from multi-modal sensing with the corresponding CSI features in latent space. This alignment allows WiFo-M2 to extract universally beneficial channel-aware features for diverse PHY actions. To bridge the sampling rate gap among different modalities, WiFo-M2 is designed as a temporal sequence-to-sequence model. It ingests a fixed historical window of sensory data and is trained to extrapolate future channel-aligned features. This enables WiFo-M2 to provide channel-aware features even when fresh sensory data is temporarily unavailable, a practical issue largely overlooked in prior work. To enhance the robustness and representation learning capability of WiFo-M2, ContraSoM incorporates modality-specific data augmentation techniques, including pixel-level transformations for images and diffusion-based feature-space augmentation specifically designed for the sparse and irregular structure of LiDAR point clouds. Our key contributions are summarized as follows:
  • $\bullet$
    We develop WiFo-M2, the first FM that empowers multi-modal environmental sensing to enhance PHY actions in a scalable, plug-and-play manner. WiFo-M2 offers a scalable solution, avoiding the cumbersome, task-specific engineering traditionally required.
  • $\bullet$
    We propose a pre-training strategy based on contrastive learning (CL), named ContraSoM. By pre-training on a large-scale multi-modal dataset, the ContraSoM strategy endows WiFo-M2 with the ability to mine OOB channel-aware features during the inter-frame perception gap.
  • $\bullet$
    We validate the capabilities of WiFo-M2 across extensive PHY actions implemented via diverse schemes. Experiments show that WiFo-M2 empowers these PHY actions without requiring fine-tuning and exhibits strong cross-scenario generalization capabilities.
II Related Work
II-A Sensing-Refined PHY Actions
Multi-modal sensing has been increasingly utilized to improve PHY actions. Existing schemes are predominantly action-specific, requiring carefully customized data processing and network designs for each application. For example, the LE-CLN scheme [7] converts raw LiDAR data into the signal propagation feature superposed range map and customizes a feature extraction neural network (NN) to transform LiDAR into OOB features beneficial for channel estimation. Similarly, the H-MVMM scheme proposed in [12] is specifically designed for precoding, with carefully crafted data processing methods for Global Positioning System (GPS), LiDAR, and RGB images, as well as customized feature extraction and fusion NNs. The MSCP channel prediction framework proposed in [5] adopts a bespoke pipeline that first aligns camera images with pilots via angle transformation and path filtering, and then employs a task-specific cross-modal Transformer to regress channel parameters. Such design principle leads to a fundamental limitation: the performance gain lacks universality and is only effective for a particular PHY action, scenario, and transceiver setup.
Another pressing issue is that current research remains concentrated on just a few PHY actions whose outputs are strongly tied to user location, such as beam prediction and handover. To date, whether environment sensing can benefit many other PHY actions is still largely unexplored. The difficulty stems from the highly diverse and often latent relationships between sensing and PHY actions. For example, temporal channel prediction (CP) relies on capturing time-varying characteristics of scatterers; uplink–downlink CP concerns spatial consistency across frequency bands. Hence, crafting a dedicated sensing pipeline and mining task-specific features becomes labor-intensive. In this context, how to obtain a universal OOB feature from multi-modal sensing and use it to boost a wide range of PHY actions becomes a critical challenge.
II-B CL-Based Foundation Model for Cross-Modal Alignment
WiFo-M2 is envisioned to extract universal OOB channel-aware features from multi-modal environment sensing. Achieving this requires a self-supervised pre-training scheme that (i) learns from large-scale unlabeled data and (ii) aligns heterogeneous modalities. CL [13], which constructs representations by learning similarities between data samples, naturally meets the above requirements. Recently, beyond vision-language domains, CL has been effectively adapted to align sensor modalities with complementary characteristics. DiffCL [14] introduces a diffusion-based graph CL module that generates denoised visual- and text-aware views, encouraging the encoder to distill robust user–item representations even under severe interaction sparsity and multi-modal noise. In medical image segmentation, CLMorph [15] integrates CL into a deformable registration framework to align an unaligned image with a reference image. By contrasting feature representations extracted from these images, it enhances the discriminative capacity of the network. Similarly, in multi-media recommendation, graph-based contrastive methods [16] are designed to disentangle and align user preferences across visual, acoustic, and textual modalities.
Building upon these advances, CL has also been introduced as an effective pre-training strategy in wireless communication community. CSI-CLIP proposed in [17] treats CSI and channel impulse response as naturally aligned pairs and performs cross-modal CL between them, enabling the model to capture generalizable channel features for improved positioning and beam management. LWLM [18] introduces position-invariant CL, which contrasts CSI samples collected from the same location under different BS configurations to learn robust and location-aware semantics specifically for localization tasks. Furthermore, in [19], a VLM-based beam prediction (BP) framework is proposed which utilizes CL to align image and LiDAR features, enforcing cross-modal consistency to enhance multi-modal fusion for beam selection. Most recently, the WMFM model [20] adopts a vanilla form of CL to align images with CSI, yet it still suffers from several limitations. Firstly, as in most prior work, it fails to resolve the intrinsic temporal mismatch between sensory data and CSI. Secondly, it utilizes solely the image modality and does not employ any data pre-processing and augmentation techniques, limiting its robustness and cross-domain generalization. Thirdly, the pre-training dataset covers only one scene and a single communication system configuration. Finally, it does not investigate how multi-modal sensing could benefit a wider range of PHY actions and remains confined to Line of sight (LoS)/non-Line of sight (NLoS) classification and localization.
III System Model
As illustrated in Fig. 2, we consider an intelligent transportation scenario, where the BS is an intelligent agent that serves $M$ single‑antenna users and senses the physical world through multi‑modal sensors. In this work, we assume that multi‑modal environment sensing includes LiDAR point clouds and RGB images, denoted by $\bm{P}_{\text{raw}}\in\mathbb{R}^{N\times 3}$ and $\bm{I}\in\mathbb{R}^{3\times H\times W}$, respectively. We assume the camera operates at $40$ fps, while the LiDAR scans at $20$ Hz.
We consider a multiple-input single-output orthogonal frequency division multiplexing (MISO-OFDM) communication system under EIW paradigm where BS is equipped with either a uniform linear array (ULA) or a uniform planar array (UPA), with the antenna size denoted by $N$ and the number of subcarriers by $K$. Let $\mathbf{H}\in\mathbb{C}^{K\times N}$ represent CSI between the BS and a certain user.
Figure 2: An overview of the considered scenario.
Figure 3: Overall framework of WiFo-M2, illustrating the two‑stage pipeline: (left) pre-training stage with ContraSoM strategy realizing SoM processing, and (right) inference stage for optimizing various PHY actions in a plug-and-play manner.
IV WiFo-M2 Framework: Sensing Processing, Feature Extraction, and Pre-training
In this section, we first analyze the fundamental challenges faced by developing WiFo-M2 to achieve scalable environment sensing-refined PHY actions for intelligent network entities. Then, we introduce the core designs: the unified environment sensing processing pipeline, the architecture and feature extraction method of the WiFo-M2, and the ContraSoM pre‑training strategy. The overall framework of WiFo‑M2 is depicted in Fig. 3.
IV-A Challenges
To realize the goal of scalable sensing-refined PHY actions, the primary obstacles arise from three key aspects: environment sensing data processing, foundation model design, and the pre-training strategy. These are detailed as follows:
(1) Absence of Inherent Cross-Modal Data Pairing: A fundamental premise for CL is the availability of aligned positive sample pairs across modalities. In wireless communication scenarios, however, a single sensing frame captures elements like vehicles and buildings can correspond to multiple potential receivers. Consequently, the first challenge lies in disentangling and constructing matching CSI positive pairs for different receivers from the same set of sensory data.
(2) Temporal Misalignment Due to Heterogeneous Sampling Rates: Wireless channels vary every few milliseconds, whereas intelligent network entities can only observe the surrounding environment at low frame rates (e.g., 1010-2020 Hz). That is, most PHY actions lack concurrent environmental observations during execution. Therefore, WiFo-M2 cannot simply rely on aligning sensory frames with coincident CSI samples. Instead, it has to effectively “fill in” the missing environmental context at the communication system’s operational tempo.
(3) Effective Multi-modal Data Augmentation Technique: Data augmentation is crucial for generating diverse, semantic-preserving views to learn robust representations. However, crafting augmentation strategies for complex, unstructured modalities like LiDAR point clouds is non-trivial. Simple point dropout or perturbation can easily corrupt geometric relationships, undermining the consistency required for CL. Effective augmentation methods should be developed to maintain the data’s fundamental spatial and semantic structure to ensure the generated views remain valid positive samples.
IV-B Unified Environment Sensing Processing Pipeline
IV-B1 Image Processing Pipeline
The image processing pipeline converts raw images containing many potential receivers into a representation deterministically linked to the CSI of a specific transmitter and receiver pair, and it comprises three core steps.
Step 1: Initial Object Detection and Tracking. The raw image data is first subjected to a pixel-level augmentation pipeline (described in detail in Section IV-D1), and the resulting image is denoted as $\tilde{\bm{I}}$. The image processing pipeline begins with object detection to identify potential communication targets (e.g., vehicles in V2X communication scenarios) from $\tilde{\bm{I}}$. We assume the BS has prior knowledge of the receiver’s initial position through initial access procedures. Then, the advanced object detection model YOLOv8 is used to detect all targets in the image and perform continuous tracking across frames.
Step 2: Receiver-Bounding Box Matching. For each detected bounding box, we calculate its angular span in the camera’s field of view. Given a bounding box with normalized coordinates $(x_{c}^{n},y_{c}^{n},w^{n},h^{n})$ where $x_{c}^{n},y_{c}^{n}$ are the normalized center coordinates and $w^{n},h^{n}$ are the normalized width and height, we first convert to pixel coordinates: $x_{c}=x_{c}^{n}\times W$, $y_{c}=y_{c}^{n}\times H$, $w=w^{n}\times W$, and $h=h^{n}\times H$, where $W$ and $H$ are the image width and height in pixels. The horizontal angular span $\theta_{\text{min}},\theta_{\text{max}}]$ of the bounding box is then computed using the camera’s intrinsic parameters:
$\theta_{\text{min}}$ $=\arctan\left(\frac{x_{\text{min}}-c_{x}}{f_{x}}\right),$ (1)
$\theta_{\text{max}}$ $=\arctan\left(\frac{x_{\text{max}}-c_{x}}{f_{x}}\right),$ (2)
where $x_{\text{min}}=x_{c}-w/2$, $x_{\text{max}}=x_{c}+w/2$, $f_{x}$ is the focal length in pixels, and $c_{x}$ is principal point offset along x-axis.
To match the receiver with the correct bounding box, we compute the receiver’s azimuth angle relative to camera. Given the camera’s world coordinates $\bm{p}_{\text{cam}}=(x_{\text{cam}},y_{\text{cam}},z_{\text{cam}})$ and yaw angle $\psi_{\text{cam}}$, and the receiver’s world coordinates $\bm{p}_{\text{rec}}=(x_{\text{rec}},y_{\text{rec}},z_{\text{rec}})$, the relative azimuth angle is:
$\theta_{\text{rec}}=\text{wrap}\left(\arctan\left(\frac{y_{\text{rec}}-y_{\text{cam}}}{x_{\text{rec}}-x_{\text{cam}}}\right)-\psi_{\text{cam}}\right),$ (3)
where $\text{wrap}(\cdot)$ ensures the angle remains within $(-\pi,\pi]$. The bounding box whose angular span contains $\theta_{\text{rec}}$ is selected as the match. If multiple bounding boxes satisfy this condition, we choose the one whose angular center $\bar{\theta}=(\theta_{\text{min}}+\theta_{\text{max}})/2$ is closest to $\theta_{\text{rec}}$.
Step 3: Receiver-Labeling via Angular Encoding. Once the matching bounding box is identified, we perform a receiver-labeling operation by replacing the original pixel content within the bounding box with a color that encodes the receiver’s angular information. Through this operation, a positive sample association between the processed image and the CSI between BS and a specific user is established, and geometric information is also directly embedded into the image for subsequent feature learning.
The angular encoding proceeds as follows. First, we clip the receiver’s azimuth angle to the range $\pi/2,\pi/2]$, i.e., $\theta_{\text{clipped}}=\max\left(-\frac{\pi}{2},\min\left(\frac{\pi}{2},\theta_{\text{rec}}\right)\right)$. Next, we normalize this angle to the range $[0,1]$ by $\theta_{\text{norm}}=(\theta_{\text{clipped}}+\pi/2)/\pi$. Finally, we map the normalized angle to a hue value in the HSV color space via $h=\lfloor 179\times\theta_{\text{norm}}\rfloor$. The corresponding RGB color $\bm{c}_{\text{RGB}}=(r,g,b)$ is obtained by converting the HSV color $(h,255,255)$ to the RGB color space. The bounding box region is then filled with this color:
$\tilde{\bm{I}}(y_{1}:y_{2},x_{1}:x_{2})=\bm{c}_{\text{RGB}},$ (4)
where $(x_{1},y_{1})$ and $(x_{2},y_{2})$ define the bounding box corners in pixel coordinates. The transformed image after this operation is denoted as $\bm{I}_{\rm labeled}$.
IV-B2 LiDAR Processing Pipeline
Given a raw point cloud $\bm{P}_{\text{raw}}\in\mathbb{R}^{N\times 3}$ consisting of $N$ points with 3D coordinates $(x,y,z)$, we first remove ground points and obtain a filtered point cloud $\bm{P}_{\text{filt}}\in\mathbb{R}^{N^{\prime}\times 3}$.
Step 1: Object Clustering and Initial Receiver Matching. To segment the filtered point cloud into distinct objects, we apply the DBSCAN clustering algorithm to $\bm{P}_{\text{filt}}$, obtaining a set of clusters $\mathcal{C}=\{C_{1},C_{2},\dots,C_{K}\}$. For each cluster $C_{k}$, we compute its centroid $\bm{c}_{k}=(x_{k},y_{k},z_{k})$. During the initial access phase, the receiver reports its global position $\bm{p}_{\text{rec}}=(x_{\text{rec}},y_{\text{rec}},z_{\text{rec}})$. The receiver‑corresponding cluster is identified as the one whose centroid is closest to $\bm{p}_{\text{rec}}$: $k^{*}=\arg\min_{k}\|\bm{c}_{k}-\bm{p}_{\text{rec}}\|$.
Step 2: Continuous Tracking and Point Labeling. Once the receiver cluster is identified in the initial frame, we employ a multi‑object tracking algorithm to maintain consistent cluster identities across consecutive frames. This ensures that the receiver vehicle is correctly associated throughout the sequence. Let the receiver cluster be denoted as $\mathcal{C}_{\text{rec}}$, and let the set of building points be defined by a geometric rule, e.g., $\mathcal{B}=\{\bm{p}_{i}\in\bm{P}_{\text{filt}}:y_{i}>y_{\text{build}}\}$. Then each point $\bm{p}_{i}=(x_{i},y_{i},z_{i})$ in $\bm{P}_{\text{filt}}$ is assigned a label $l_{i}\in\{-1,0,1\}$ according to:
$l_{i}=\begin{cases}1,&\text{if }\bm{p}_{i}\in\mathcal{C}_{\text{rec}}\\ -1,&\text{if }\bm{p}_{i}\in\mathcal{B}\setminus\mathcal{C}_{\text{rec}}\\ 0,&\text{otherwise}\end{cases}.$ (5)
Finally, the labeled point cloud $\bm{P}_{\text{labeled}}\in\mathbb{R}^{N^{\prime}\times 4}$ is then constructed by appending the label column to $\bm{P}_{\text{filt}}$, i.e., $\bm{P}_{\text{labeled}}=[\bm{P}_{\text{filt}},\bm{l}]$, where $\bm{l}=[l_{1},l_{2},\dots,l_{N^{\prime}}]^{T}$.
IV-C Network Architecture and Working Flow of WiFo-M2
The WiFo-M2 consists of two modality-specific backbones: WiFo-M2-Img built upon ResNet-34 [21], and WiFo-M2-LiDAR built upon PointNet [22]. Essentially, these two backbones act as modality-specific feature extractors. Their architectures are depicted in Fig. 4. To capture temporal dependencies in sequential sensing data, WiFo-M2-Img and WiFo-M2-LiDAR are augmented with Gated Recurrent Units (GRU) for sequential processing.
IV-C1 CSI Feature Extraction
For CSI feature extraction, we employ the pre-trained WiFo model [23] as a powerful space-temporal-frequency (STF) feature extractor. Given an input CSI data $\mathbf{H}_{t}\in\mathbb{C}^{K\times N}$ at timestamp $t$, the WiFo model processes it through its transformer-based encoder. The output is a fixed-dimensional feature vector $\mathbf{z}^{\text{C}}_{t}\in\mathbb{R}^{d}$.
IV-C2 Image and LiDAR Feature Extraction
Both image and LiDAR modalities follow the same feature extraction procedure comprising three stages:
Figure 4: Network architecture of WiFo-M2-Img/LiDAR.
Stage 1: Spatial Feature Extraction. Let $T$ denote the timestamp of the most recent sensory data. Define $\Delta t_{\rm I}$ and $\Delta t_{\rm L}$ as the inter-frame intervals of the image and LiDAR streams, respectively, and $(n_{\rm I}+1)$ and $(n_{\rm L}+1)$ as the sequence lengths of the LiDAR and image inputs. For image modality, each frame $\bm{I}^{t}_{\rm labeled}$ passes through the ResNet-34 backbone of the WiFo-M2-Img model to obtain frame-level features. For LiDAR modality, each point cloud frame $\bm{P}^{t}_{\rm labeled}$ is processed by the PointNet backbone [22] to extract frame-level features. The extraction process can be formally expressed as:
$\bm{u}_{t}^{\text{I}}$ $=f_{\text{ResNet}}\left(\bm{I}^{t}_{\rm labeled}\right)\in\mathbb{R}^{d},\ t=T-n_{\rm I}\Delta t_{\rm I},\ldots,T$ (6)
$\bm{u}_{t}^{\text{L}}$ $=f_{\text{PointNet}}\left(\bm{P}^{t}_{\rm labeled}\right)\in\mathbb{R}^{d},\ t=T-n_{\rm L}\Delta t_{\rm L},\ldots,T$ (7)
where $f_{\text{ResNet}}$ and $f_{\text{PointNet}}$ denote the ResNet-34 and PointNet backbone networks respectively. We denote the frame-level features for image and LiDAR within a time window $T-n_{\rm I}\Delta t_{\rm I},T]$ or $T-n_{\rm L}\Delta t_{\rm L},T]$ as $\bm{u}_{t}^{\text{I}}\in\mathbb{R}^{d}$ and $\bm{u}_{t}^{\text{L}}\in\mathbb{R}^{d}$, respectively.
Stage 2: Temporal Feature Modeling. To capture temporal dependencies, both modalities employ the GRU network that learns the underlying dynamics of the environment. The GRU processes the sequences of intermediate frame-level features $\{\bm{u}^{\rm x}_{t}\}_{t=T-n_{\rm x}\Delta t_{\rm x}}^{T},{\rm x}\in\{\rm L,\rm I\}$ to produce a sequence of hidden states, which are then projected to the final feature space. Let $\Delta\tau$ be the PHY action execution interval. The resulting projected feature sequence corresponds to the prediction window $T,T+\Delta\tau,\cdots,T+\Delta t_{\rm x}]$, i.e., from the current sensing timestamp $T$ up to the moment when the next sensing frame is captured. This operation can be expressed as:
$\bm{v}^{\rm x}_{\tau}\$$=\phi_{\rm x}\Bigl(\mathcal{G}_{\rm x}\left(\{\bm{u}^{\rm x}_{t}\}_{t=T-n_{\rm x}\Delta t_{\rm x}}^{T}\right)\Bigr),\tau=T,\cdots,T+\Delta t_{\rm x}$ (8)
where $\mathcal{G}_{\rm x}$ denotes the GRU network for modality $\rm x$, $\phi_{\rm x}$ represents the projection layer, and $\bm{v}_{t}^{\rm x}\in\mathbb{R}^{d}$ is the final frame-level feature at time step $t$ for modality $\rm x$. The sequence of final features $\{\bm{v}^{\rm x}_{\tau}\}_{\tau=T}^{T+\Delta t_{\rm x}}$ is then used for subsequent temporal alignment and contrastive learning.
Figure 5: Temporal feature extrapolation in WiFo-M2-Img.
Stage 3: Multi-Timestamp Feature Generation and Alignment. A key design of WiFo-M2’s architecture is that the GRU generates channel-aware features for multiple future timestamps. During pre-training, the sequence of final features $\{\bm{v}^{\rm x}_{\tau}\}_{\tau=T}^{T+\Delta t_{\rm x}},{\rm x}\in\{\rm L,\rm I\}$ generated by WiFo-M2 are individually aligned with the CSI feature at the corresponding timestamp. For a given timestamp $\tau$, a certain sensing feature $\bm{v}^{\rm{x}}_{\tau}$ is selected and paired with the CSI feature $\bm{z}^{\text{C}}_{\tau}$ from the same timestamp to form a positive pair for CL (detailed in the next section). As shown in Fig. 5, this design enables WiFo-M2 to learn to generate temporally aligned features at specific future timestamps, which can then be directly compared with CSI features from the corresponding time in CL objective, ensuring precise cross-modal temporal alignment.
IV-D Pre-training Strategy for WiFo-M2: ContraSoM
This section presents the proposed ContraSoM pre-training strategy in detail, covering the modality-specific data-augmentation techniques together with the overall contrastive learning pipeline. By aligning multi-modal environment sensing and wireless channel in the latent space, ContraSoM essentially performs the SoM processing outlined in [4].
IV-D1 Modality-Specific Data Augmentation
Image data possess a regular grid structure and dense pixel representation, making them amenable to direct augmentation in the pixel space. We apply a series of stochastic transformations to generate robust augmented views for contrastive learning. Given an input image $\bm{I}$, the augmented view $\tilde{\bm{I}}$ is obtained via: $\tilde{\bm{I}}=\mathcal{T}_{\text{aug}}(\bm{I})$. $\mathcal{T}_{\text{aug}}$ is a combination of several of the following operations, randomly applied to each original image: First, color jitter randomly perturbs brightness, contrast, saturation, and hue, mimicking natural variations in illumination: $\bm{I}_{\text{jit}}=\bm{I}\odot\mathbf{c}_{\text{scale}}+\mathbf{c}_{\text{shift}}$, where $\odot$ denotes element-wise multiplication, $\mathbf{c}_{\text{scale}}\sim\mathcal{U}(0.7,1.3)$ and $\mathbf{c}_{\text{shift}}\sim\mathcal{U}(-0.3,0.3)$ for each channel. Second, Gaussian Blur is applied with probability $p=0.5$ to simulate lens defocus or motion blur: $\bm{I}_{\text{blur}}=\bm{I}*G_{\sigma}$ with $\sigma\sim\mathcal{U}(0.1,2.0)$, where * denotes the convolution operation and $G_{\sigma}$ is a Gaussian kernel of size $15\times 15$. Next, random erasing is performed with probability $p=0.3$, which randomly selects a rectangular region $\mathcal{R}$ within the image and replaces the pixel values in that region with random noise, emulating occlusions or sensor dropout. This operation is defined as:
$\bm{I}_{\text{erase}}(x,y)=\begin{cases}\eta,&\text{if }(x,y)\in\mathcal{R}\\ \bm{I}(x,y),&\text{otherwise}\end{cases},$ (9)
where $\eta\sim\mathcal{U}(0,1)$, and the region $\mathcal{R}$ has an area ratio $\sim\mathcal{U}(0.01,0.05)$ and an aspect ratio $\sim\mathcal{U}(0.3,3.3)$. Finally, normalization is applied using ImageNet statistics: $\bm{I}_{\text{norm}}=(\bm{I}-\bm{\mu})/\bm{\sigma}$ with $\bm{\mu}=[0.485,0.456,0.406]$ and $\bm{\sigma}=[0.229,0.224,0.225]$. This pipeline encourages the model to learn invariant representations under appearance variations, occlusion, and noise.
LiDAR point clouds are sparse and geometrically structured. Applying noise directly in the raw data level would easily disrupt semantic relationships. Therefore, we propose to perform augmentation in the feature space. We adopt a denoising diffusion probabilistic model (DDPM) framework to generate augmented features in a controlled manner.
The augmentation process begins with a forward noising step. Taking the final LiDAR feature at a specific time step $i$, denoted as $\bm{v}^{\rm L}_{i}$, as an example, we gradually inject Gaussian noise into this feature over a series of $M$ discrete steps. This forward noising process is defined as a Markov chain:
$q(\bm{v}_{m}\mid\bm{v}^{\rm L}_{m-1})=\mathcal{N}\left(\bm{v}^{\rm L}_{m};\sqrt{1-\beta_{m}}\,\bm{v}^{\rm L}_{m-1},\beta_{m}\mathbf{I}\right),$ (10)
where $\beta_{m}$ follows a linearly increasing schedule from $\beta_{\min}=5\times 10^{-6}$ to $\beta_{\max}=2\times 10^{-3}$, $m=1,\dots,M$. Due to the additive property of Gaussian noise, we can directly sample the feature at any step $m$:
$\tilde{\bm{v}}^{\rm L}_{m}=\sqrt{\bar{\alpha}_{m}}\,\bm{v}^{\rm L}_{i}+\sqrt{1-\bar{\alpha}_{m}}\,\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})$ (11)
with $\alpha_{m}=1-\beta_{m}$ and $\bar{\alpha}_{m}=\prod_{s=1}^{m}\alpha_{s}$.
For efficient feature generation in the reverse process, we use a deterministic skip-step reverse diffusion sampler. Starting from a noisy feature $\tilde{\bm{v}}^{\rm L}_{m}$, we update with stride $\Delta m=\max(\lfloor m/u\rfloor,1)$ (where $u=8$) as
$\tilde{\bm{v}}^{\rm L}_{m-\Delta m}=$ $\sqrt{\bar{\alpha}_{m-\Delta m}}\left(\frac{\tilde{\bm{v}}^{\rm L}_{m}-\sqrt{1-\bar{\alpha}_{m}}\,\epsilon_{\theta}(\tilde{\bm{v}}^{\rm L}_{m},m)}{\sqrt{\bar{\alpha}_{m}}}\right)$
$+\sqrt{1-\bar{\alpha}_{m-\Delta m}}\,\epsilon_{\theta}(\tilde{\bm{v}}^{\rm L}_{m},m).$ (12)
After $U=24$ reverse steps, we obtain the augmented feature $\tilde{\bm{v}}^{\rm L}_{i}$. To enhance diversity, we generate two distinct augmented views $\tilde{\bm{v}}_{i}^{\rm L(1)}$ and $\tilde{\bm{v}}_{i}^{\rm L(2)}$ by initializing from different noise levels and noise draws, then applying the same reverse update:
$\tilde{\bm{v}}^{\rm L}_{m_{j}}=\sqrt{\bar{\alpha}_{m_{j}}}\,\bm{v}^{\rm L}_{i}+\sqrt{1-\bar{\alpha}_{m_{j}}}\,\epsilon_{j},\quad j\in\{1,2\},$ (13)
where $m_{1},m_{2}\sim\mathcal{U}\{200,400\}$ and $\epsilon_{1},\epsilon_{2}\sim\mathcal{N}(0,\mathbf{I})$. Then, run the above deterministic skip-step reverse updates for $U$ steps to obtain
$\tilde{\bm{v}}_{i}^{\rm L(1)}=\mathcal{G}_{\text{rev}}(\tilde{\bm{v}}^{\rm L}_{m_{1}}),\qquad\tilde{\bm{v}}_{i}^{\rm L(2)}=\mathcal{G}_{\text{rev}}(\tilde{\bm{v}}^{\rm L}_{m_{2}}),$
where $\mathcal{G}_{\text{rev}}(\cdot)$ denotes the $U$-step deterministic reverse sampler defined above. The two independent pairs $(m_{1},\epsilon_{1})$ and $(m_{2},\epsilon_{2})$ yield two complementary augmented views.
IV-D2 Contrastive Learning Objectives in WiFo-M2
We design two distinct CL frameworks for image-CSI alignment in WiFo-M2-Img and for LiDAR-CSI alignment in WiFo-M2-LiDAR.
For WiFo-M2-Img, we employ a symmetric contrastive learning objective between the extracted image feature $\bm{v}^{\text{I}}$ and the CSI feature $\mathbf{z}^{\text{C}}$. To ensure balanced alignment between the two modalities, we define a symmetric InfoNCE loss:
$\mathcal{L}^{\text{I}}=\mathcal{L}_{\text{sym}}(\bm{v}^{\text{I}},\mathbf{z}^{\text{C}}),$ (14)
where the symmetric InfoNCE loss for two feature sets $\mathbf{Z}_{a}$ and $\mathbf{Z}_{b}$ is:
$\mathcal{L}_{\text{sym}}(\mathbf{Z}_{a},\mathbf{Z}_{b})=\frac{1}{2}\left[\mathcal{L}_{\text{InfoN}}(\mathbf{Z}_{a},\mathbf{Z}_{b})+\mathcal{L}_{\text{InfoN}}(\mathbf{Z}_{b},\mathbf{Z}_{a})\right].$ (15)
The standard InfoNCE loss for a batch of $N$ feature pairs is defined as:
$\mathcal{L}_{\text{InfoN}}(\mathbf{Z}_{a},\mathbf{Z}_{b})=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp\left(s(\mathbf{z}_{a}^{(i)},\mathbf{z}_{b}^{(i)})/\tau_{\rm e}\right)}{\sum_{j=1}^{N}\exp\left(s(\mathbf{z}_{a}^{(i)},\mathbf{z}_{b}^{(j)})/\tau_{\rm e}\right)},$ (16)
where $\mathbf{z}_{a}^{(i)}$ and $\mathbf{z}_{b}^{(i)}$ form a positive pair, $s(\cdot,\cdot)$ denotes the cosine similarity function, and $\tau_{\rm e}$ is a temperature hyperparameter. The symmetric formulation ensures bidirectional alignment between image and CSI feature spaces.
For the LiDAR modality, we incorporate both the original LiDAR feature $\bm{v}^{\rm L}$ and the two diffusion-augmented features $\tilde{\bm{v}}^{\rm L(1)}$, $\tilde{\bm{v}}^{\rm L(2)}$ into the CL framework. The overall objective consists of two components: a contrastive loss $\mathcal{L}^{\rm L}_{\text{contra}}$ and a diffusion training loss $\mathcal{L}_{\text{diff}}$. Specifically, we formulate a triplet contrastive loss to enforce consistency between three pairs: the original LiDAR features and CSI features, and the augmented LiDAR features and CSI features:
$\mathcal{L}_{\text{contra}}=$ $\lambda_{1}\cdot\mathcal{L}_{\text{sym}}(\bm{v}^{\rm L},\mathbf{z}^{\text{C}})+\lambda_{2}\cdot\mathcal{L}_{\text{sym}}(\tilde{\bm{v}}^{\rm L(1)},\mathbf{z}^{\text{C}})$
$+\lambda_{3}\cdot\mathcal{L}_{\text{sym}}(\tilde{\bm{v}}^{\rm L(2)},\mathbf{z}^{\text{C}}).$ (17)
In parallel to the above contrastive objective, we incorporate a diffusion training loss to optimize the sampler $\epsilon_{\theta}(\cdot)$. This component follows the standard DDPM objective, minimizing the mean-squared error between the predicted and actual noise:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{m,\bm{v}^{\rm L},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\tilde{\bm{v}}^{\rm L}_{m},m)\|^{2}\right],$ (18)
with $m\sim\mathcal{U}\{1,\dots,T\}$ and $\epsilon\sim\mathcal{N}(0,\mathbf{I})$.
The complete training objective for WiFo-M2-LiDAR thus combines both losses with a balancing weight:
$\mathcal{L}^{\text{L}}=\mathcal{L}_{\text{contra}}+\lambda_{\text{diff}}\cdot\mathcal{L}_{\text{diff}},$ (19)
where $\lambda_{\text{diff}}=0.3$ in our implementation.
V Experiments
In this section, we begin by detailing the construction of the pre-training and testing datasets. Next, we provide a detailed description of the network architecture and training hyper-parameters. Finally, we comprehensively evaluate and analyze the performance of WiFo-M2.
V-A Datasets
TABLE I: Overview of Pre-training Datasets.
Dataset Type Data Source Scenario Carrier Frequency (GHz) Dataset ID BS Antenna Configuration Bandwidth (GHz) Number of Subcarriers Samples
Pre-training Dataset M3SC [26] Intersection traffic $28$ I1 $32\times 1$ $0.2$ $512$ $7,048$
I2 $64\times 1$ 11 $1024$ $13,888$
I3 $16\times 8$ $0.8$ $512$ $7,088$
I4 $16\times 8$ $0.4$ $512$ $7,168$
I5 $8\times 8$ $0.8$ $1024$ $14,048$
I6 $8\times 4$ 11 $256$ $3,484$
SynthSoM [27] Dense building $4.95$ B1 $16\times 16$ $0.05$ $256$ $9,080$
B2 $16\times 8$ $0.05$ $512$ $16,784$
B3 $8\times 8$ $0.05$ $512$ $14,816$
B4 $8\times 4$ $0.1$ $512$ $18,824$
B5 $4\times 4$ $0.1$ $1024$ $38,064$
TABLE II: Overview of Testing Datasets.
Dataset Type Data Source Scenario Carrier Frequency (GHz) Dataset ID BS Antenna Configuration Bandwidth (GHz) Number of Subcarriers Samples
Testing Set (Seen Scenarios) SynthSoM [27] M3SC [26] Intersection traffic & Dense building $28$, $4.95$ L1 $16\times 8$ $0.8$ $0.05$ $512$ $809$
L2 $4\times 4$ $0.1$ $1024$ $547$
L3 $16\times 16$ $0.05$ $256$ $300$
L4 $8\times 4$ 11 $256$ $229$
L5 $32\times 1$ $8\times 4$ $0.2$ $0.1$ $512$ $725$
L6 $64\times 1$ $8\times 8$ 11 $0.8$ $64$ $404$
L7 $8\times 8$ $0.05$ $512$ $437$
Testing Set (Unseen Scenarios) SynthSoM-Twin [29] Campus $5.9$ C1 $64\times 1$ $0.02$ $64$ $480$
DeepSense-6G [31] (Measured data) Scenario 31 $60$ S1 $16\times 1$ - - $387$
ViWi [30] dist_cam $60$ V1 $64\times 1$ $0.5$ $256$ $1302$
We construct a multi-modal pre-training dataset that includes LiDAR, RGB images, and CSI data. The CSI data is collected from the links between the BS and the users. In all datasets, the users are equipped with a single antenna, while the BS employs either a UPA or ULA configuration, with adjacent antenna spacing set to half the wavelength at the central frequency. The data is sourced from the intersection traffic scenario in the M3SC dataset [26] and the dense building scenario in the SynthSoM dataset [27]. Various communication system configurations are considered, covering a wide range of subcarrier numbers, bandwidths, and antenna sizes. The detailed simulation configurations of the pre-training datasets are presented in Table I.
We also construct a testing set, with detailed simulation configurations shown in Table II. The testing set consists of two parts: $7$ datasets for seen scenarios and $3$ datasets for unseen scenarios. First, to validate the in-distribution performance of WiFo-M2, we construct datasets L1 to L7, which are selected from the same links as the pre-training datasets. Among these, datasets L1, L5, and L6 are formed by merging several links with the same number of antennas and subcarriers but smaller data volumes. To test the cross-scenario generalization ability of WiFo-M2, we construct datasets C1, S1, and V1 based on three new scenarios not included in the pre-training datasets, namely, the campus scenario in the SynthSoM-Twin dataset [29], the “dist_cam” scenario in the ViWi dataset [30], and the Scenario 31 in the DeepSense-6G dataset [31], respectively. Notably, these three datasets operate at entirely new carrier frequencies, and their images and LiDAR point clouds exhibit visual appearances and spatial layouts unseen in the pre-training data. This pronounced gap between training and evaluation domains highlights the strong universality and generalization capability of WiFo‑M2. Fig. 6 provides an overview of the multi-modal environment sensory data examples from all the aforementioned scenarios.
Figure 6: Overview of multi-modal sensory data from all datasets.
V-B Network and Pre-training Settings
The WiFo-M2-LiDAR is trained for $100$ epochs using a batch size of $1024$. The optimization employs the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and a weight decay of $5\times 10^{-4}$. A cosine annealing learning rate scheduler with a warm-up phase is utilized: the learning rate linearly increases from zero to the target learning rate over the first $10$ epochs, then decays following a cosine schedule to zero by the final epoch. The contrastive learning objective is optimized with an initial temperature parameter $\tau_{\rm e}=0.1$ and the output feature dimension $d$ set to $512$. We set the loss weights in Eq. (IV-D2) to $\lambda_{1}=1$, $\lambda_{2}=0.25$, and $\lambda_{3}=0.25$. The WiFo-M2-Img shares the same hyperparameter settings as WiFo-M2-LiDAR, except that the learning rate is set to $5\times 10^{-4}$.
V-C Performance Evaluation
In this work, we select four representative PHY actions: beam prediction, pilot-based channel estimation, channel interpolation and channel prediction, which together span the whole CSI pipeline from acquisition to utilization and underpin reliable network connectivity, providing a comprehensive test bed for our framework.
The testing dataset is further divided into a training set and a validation set with a ratio of $4:1$, referred to as the T-T dataset and the T-V dataset, respectively. Note that in all subsequent experiments, the existing schemes are trained on T-T dataset and their performance is evaluated on T-V dataset. During this process, the weight of the WiFo-M2 remains frozen, and only the existing schemes and the adapter module are trained.
V-C1 PHY Action 1-Beam Prediction
  • Problem Description: BP refers to directly predicting the optimal beamforming vector $\mathbf{f}^{*}\in\mathcal{F}$ from a pre-defined codebook $\mathcal{F}=\{\mathbf{f}_{q}\}_{q=1}^{Q}$ using multi-modal sensing. Let $\mathbf{h}\in\mathbb{C}^{1\times N}$ be the CSI of a certain subcarrier. The optimal vector $\mathbf{f}^{*}=\mathop{\arg\max}_{m\in\{1,2,\cdots,Q\}}|\mathbf{h}\mathbf{f}_{m}|^{2}$ results in the maximum effective channel gain and should be chosen for transmission. The BP action can be formulated as: $\mathbf{f}^{*}=\mathcal{Q}_{\rm BP}(\mathbf{M})$, where $\mathcal{Q}_{\rm BP}(\cdot)$ represents a certain BP scheme and $\mathbf{M}$ is the multi‑modal sensory data.
    Unlike the other three PHY actions, BP entirely relies on multi-modal sensing. Therefore, the WiFo-M2 model is directly connected to a $2$-layer multilayer perceptron (MLP) network to output the optimal beam index. During fine-tuning on the T-T dataset, only the MLP is trained. Under this setting, the BP performance intuitively reflects the effectiveness of the OOB channel-aware features.
  • Existing BP schemes: We selected two existing BP schemes, namely Vision-BP [10] and MM-BP [28]. Vision-BP [10] detects transmitters with the YOLOv3 and then feeds the bounding-box centers to a NN to select the beam index. MM-BP [28] is a multi‑modal fusion based scheme that extracts features from images, radar, and LiDAR.
Figure 7: The Top-$k$ accuracies ($k\in\{1,3,5\}$) of the WiFo-M2-based BP scheme, Vision-BP scheme, and MM-BP scheme.
  • Experiment Settings: The training hyperparameters for Vision-BP and MM-BP schemes are kept identical to those used in [10] and [28]. The proposed WiFo-M2-based BP scheme is trained with a batch size of $32$ and an initial learning rate of $2\times 10^{-4}$.
  • WiFo-M2’s Performance: We evaluate the performance of BP using the top-$k$ accuracy metric. This metric measures the percentage of test samples for which the ground-truth optimal beam is included in the model’s $k$ highest-ranked predictions. In our experiments, we report results for $k=1,3,5$ to comprehensively assess the prediction capability.
    As shown in Fig. 7, across all $7$ datasets, WiFo-M2 models exhibit highly competitive prediction accuracy even though the backbones are frozen. The WiFo-M2-Img scheme attains noticeably higher Top-1 prediction accuracy than Vision-BP scheme on L4, L5 and L6, whereas on the remaining four datasets its performance is only slightly lower than that of Vision-BP. This demonstrates that the WiFo-M2-Img backbone acquires channel-aware knowledge extracted during pre-training. The effect is more pronounced on LiDAR-based BP: WiFo-M2-LiDAR substantially surpasses the fully-trained MM-BP-LiDAR baseline on every dataset, which indicates that the pre-trained WiFo-M2-LiDAR captures spatial patterns that are highly predictive of the optimal beam. When both modalities are available, WiFo-M2 delivers virtually the same performance as the MM-BP scheme, with a mean Top-1 gap of only 2%. In summary, the BP results confirm that the ContraSoM pre-training aligns visual and LiDAR features with the wireless channel domain. Consequently, high-quality beam selection can be achieved with minimal retraining overhead.
V-C2 PHY Action 2-Channel Estimation
  • Problem Description: CE action refers to the process in which the BS estimates the channel $\mathbf{h}$ at the pilot-placed subcarriers. We assume that the BS adopts the hybrid (analog/digital) structure and is equipped with $N_{\rm{RF}}$ RF chains. The digital and analog precoding matrices are represented by $\mathbf{F}_{\rm{B}}\in\mathbb{C}^{N_{\rm{RF}}\times N_{\rm{RF}}}$ and $\mathbf{F}_{\rm{R}}\in\mathbb{C}^{N\times N_{\rm{RF}}}$ with $\|\mathbf{F}_{\rm{R}}\mathbf{F}_{\rm{B}}\|_{\rm{F}}^{2}=N_{\rm{RF}}$. $\mathbf{F}_{\rm{B}}$ is set as an identity matrix $\mathbf{I}_{N_{\rm{RF}}}$ during CE. The quantization bit number of the phase shifters adopted at BS is set to $B=3$. Each element of $\mathbf{F}_{\rm{R}}$ is randomly chosen from the set $\mathcal{A}_{Q}=\bigl\{\,e^{j\frac{2\pi q}{Q}}\,\bigm|\,q=1,\dots,Q\bigr\}$, where $Q\triangleq 2^{B}=8$. During CE, the single-antenna user transmits known pilots $\mathbf{x}\in\mathbb{C}^{1\times T}$ over $T$ successive time slots. The received signal sequence at BS $\mathbf{Y}\in\mathbb{C}^{N_{\rm RF}\times T}$ can then be expressed as:
    $\mathbf{Y}=\mathbf{F}_{\rm{B}}^{\rm H}\mathbf{F}_{\rm{R}}^{\rm H}\mathbf{h}^{\rm H}\mathbf{x}+\mathbf{F}_{\rm{B}}^{\rm H}\mathbf{F}_{\rm{R}}^{\rm H}\mathbf{N},$ (20)
    where $\mathbf{N}\in\mathbb{C}^{N\times T}$ denotes the additive white Gaussian noise with i.i.d. entries of zero mean and variance $\sigma^{2}$. Upon receiving the pilots, BS employs a certain CE network to estimate the CSI: $\hat{\mathbf{h}}=\mathcal{Q}_{\text{CE}}\bigl(\mathbf{Y},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CE}(\cdot)$ represents a certain CE scheme and $\mathcal{G}(\cdot)$ denotes the WiFo‑M2 model.
  • Existing CE schemes: We selected two existing CE schemes, namely CENN [32] and FCDAMP [33]. CENN [32] predicts the beam-space channel magnitudes and reconstructs the channel from the dominant coefficients, whereas FCDAMP [33] inserts a fully convolutional denoising network into a learned approximate message passing loop to iteratively remove noise and recover the channel.
  • Experiment Settings: The number of time slots $T$ is set to $8$ for the CENN scheme and $16$ for the FCDAMP scheme, while the number of RF chains $N_{\rm RF}$ is $4$ for all test datasets and schemes. The learning rate for all schemes is set to $1\times 10^{-4}$, and the batch size is $32$.
  • WiFo-M2’s Performance: The normalized mean square error (NMSE) is used as the estimation performance metric given by ${\rm NMSE}={{\|\hat{\mathbf{H}}-\mathbf{H}\|}_{\rm{F}}^{2}}/{{\|\mathbf{H}\|}_{\rm{F}}^{2}}$. We evaluate the channel estimation performance of CENN and FCDAMP schemes under varying SNR conditions. As shown in Fig. 8, the WiFo-M2-enhanced variants consistently outperform the original CENN and FCDAMP models across most datasets and SNRs, demonstrating the general utility of the OOB channel-aware features. Furthermore, multi-modal fusion (WiFo-M2) typically yields the most robust and significant gains, with LiDAR-only and image-only enhancements providing comparable improvements in most scenarios. On the L3 dataset with a large $256$-antenna array and limited pilots, the performance gains are less pronounced and sometimes irregular. This aligns with the expectation that under extremely sparse measurements and high-resolution array requirements, the coarse prior provided by multi-modal sensing becomes less effective in refining the estimation.
Figure 8: CE Performance of CENN and FCDAMP with WiFo-M2, WiFo-M2-Img, and WiFo-M2-LiDAR.
V-C3 PHY Action 3-Channel Interpolation
  • Problem Description: During the CE stage, pilots are placed only on a subset of subcarriers, and a subset of BS’s antenna elements is activated [^1] In CI action, we assume that a subset of antennas are provided with coarse channel estimates, while the CI action are required to recover the full CSI across all antennas. Although this setup does not reflect a practical system configuration, it serves as an effective testbed for evaluating the evaluated CI schemes’ ability to perform spatial super‑resolution reconstruction.. Upon receiving pilots, BS estimates the channel at pilot positions using the Least Squares (LS) algorithm. All the available channel estimates form the low-resolution (LR) channel observation, denoted by $\hat{\mathbf{H}}^{\rm LR}\in\mathbb{C}^{K_{\rm p}\times N_{\rm p}}$. Specifically, if pilots are placed every $k$ subcarrier and every $x$ antenna element, the LR observation can be expressed as:
    $\hat{\mathbf{H}}^{\rm LR}[i,j]=\mathbf{H}[1:k:K,1:x:N]+\mathbf{N}[i,j],$ (21)
    where $\mathbf{N}$ denotes the estimation noise. Then, BS utilizes a certain CI scheme to recover the high-resolution (HR) channel image $\hat{\mathbf{H}}\in\mathbb{C}^{K\times N}$. The entire process can be formulated as: $\hat{\mathbf{H}}=\mathcal{Q}_{\rm CI}(\hat{\mathbf{H}}^{\rm LR},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CI}(\cdot)$ represents a certain CI scheme.
Figure 9: Performance gain brought by WiFo-M2, WiFo-M2-Img, and WiFo-M2-LiDAR on LPCCNet and SR-CI.
  • Existing CI schemes: We selected two existing CI schemes, namely SR-CI [35] and LPCCNet [34]. SR-CI [35] obtains a low-precision and low-resolution channel matrix via LS and then feeds the linearly interpolated result into a super-resolution network to enhance the accuracy of CI. LPCCNet [34] is a Laplacian pyramid‑based channel completion network that progressively reconstructs the high‑resolution channel image from a low‑resolution input through a stepwise up‑sampling architecture.
  • Experiment Settings: We assume that the pilot placement interval in the subcarrier dimension and the sampling ratios for antenna dimensions are set to $k=K/8$ and $x=N/8$, respectively. That is, coarse channel estimates are available on only $1/8$ of the subcarriers and $1/8$ of the antennas, respectively. For network hyper-parameters, the SR-CI scheme uses a learning rate of $1\times 10^{-2}$, the LPCCNet scheme uses $1\times 10^{-4}$. The batch size for both schemes is $32$.
  • WiFo-M2’s Performance: The performance gains of WiFo-M2 models over LPCCNet and SR-CI schemes are illustrated in Fig. 9. Among the three variants, WiFo-M2 consistently achieves the most stable and significant improvements across different datasets. This demonstrates the effectiveness of multi-modal fusion in enhancing CI accuracy. Note that on the L3 dataset the WiFo-M2 models introduce a marginal performance degradation. A likely reason is that the L3 dataset configures a large antenna array, which translates into a very fine angular-resolution requirement for CI. In such a regime, the spatial side information from multi-modal sensing remain too coarse or imperfectly aligned with the complex channel, so the supplementary features cannot improve the high-precision estimates demanded by LPCCNet. Moreover, the performance gains brought by WiFo-M2 are especially pronounced on the L4 and L6 datasets. L4 employs a BS with a small antenna array, while L6 is configured with relatively fewer subcarriers. Under these reduced spatial- or frequency-resolution settings, the OOB features provided by WiFo-M2 are just sufficient to recover the correspondingly lower-dimensional CSI, leading to the sharp performance boost.
V-C4 PHY Action 4-Channel Prediction
  • Problem Description: In frequency division duplex (FDD) systems, where uplink-downlink reciprocity does not hold, the BS relies on uplink channel feedback to obtain downlink CSI. However, the feedback overhead scales linearly with the number of antennas and devices, making it infeasible for massive MIMO systems. Given that uplink and downlink transmissions experience the same physical environment, it is reasonable to utilize the uplink CSI measured at the BS to predict the downlink CSI. Similarly to [36, 37], the first half of the subcarriers is considered as the uplink channel, while the remaining half is treated as the downlink channel. Let $K_{\rm in}$ be the number of subcarriers in the first half and $K_{\rm out}$ represent the number of subcarriers in the second half. Then, the uplink channel $\mathbf{H}_{\rm d}$ and downlink channel $\mathbf{H}_{\rm u}$ can be expressed as:
    $\mathbf{H}_{\rm d}=\mathbf{H}[1:K_{\rm in},:],\mathbf{H}_{\rm u}=\mathbf{H}[K_{\rm in}+1:K_{\rm in}+K_{\rm out},:].$ (22)
    The CP action can be formulated as: $\mathbf{H}_{\rm u}=\mathcal{Q}_{\rm CP}({\mathbf{H}}_{\rm d},\mathcal{G}(\mathbf{M}))$, where $\mathcal{Q}_{\rm CP}(\cdot)$ represents a certain CP scheme.
  • Existing CP schemes: We selected two CP schemes, namely WiFo [23] and Transformer-based scheme [38]. WiFo [23] is a wireless FM designed for time/frequency-domain CP, which is pretrained via several self-supervised training tasks. The transformer-based scheme [38] is a parallel CP scheme.
TABLE III: CP performance of the WiFo and Transformer-based models, along with their variations enhanced by different WiFo-M2 models. The left and right sides of the slash represent the WiFo and Transformer-based schemes, respectively.
Schemes Datasets L1 L2 L3 L4 L5 L6 L7
CP baselines 5.88-5.88/0.60-0.60 2.41-2.41/0.91-0.91 4.51-4.51/1.38-1.38 0.93-0.93/2.74-2.74 2.36-2.36/0 4.61-4.61/1.15-1.15 2.78-2.78/1.22-1.22
CP baselines + WiFo-M2-Img 5.84-5.84/0.70-0.70 2.75-2.75/0 5.19-5.19/2.05-2.05 1.15-1.15/3.25-3.25 2.83-2.83/$1.21$ 4.73-4.73/1.13-1.13 3.05-3.05/1.21-1.21
CP baselines + WiFo-M2-LiDAR 5.92-5.92/0.69-0.69 2.68-2.68/1.43-1.43 5.07-5.07/1.90-1.90 1.09-1.09/3.12-3.12 3.41-3.41/0 5.04-5.04/1.32-1.32 2.94-2.94/1.07-1.07
CP baselines + WiFo-M2 6.02-6.02/0.76-0.76 2.12-2.12/1.12-1.12 5.27-5.27/2.25-2.25 1.27-1.27/3.64-3.64 3.61-3.61/0 5.10-5.10/1.28-1.28 3.30-3.30/1.22-1.22
  • Experiment Settings: The learning rates for the Transformer-based scheme and WiFo are set to $1\times 10^{-3}$ and $1\times 10^{-5}$, respectively. The batch size is set to $32$. The network input $\mathbf{H}_{\rm d}$ is assumed to be noisy CSI estimates with an NMSE of 20-20dB.
  • WiFo-M2’s Performance: The performance enhancements provided by WiFo-M2 models on top of baseline WiFo and Transformer-based schemes are summarized in Table III. The NMSE is also used as performance metric. The WiFo-M2 model leveraging multi-modality demonstrates its superiority by delivering highest gains for both CP baselines across most datasets, underscoring the benefit of cross-modal feature fusion. Consistent with observations in CI action, performance gains are particularly significant on datasets L4 and L6. Notably, WiFo-M2’s effectiveness is inherently bounded by the convergence of the baseline scheme itself. For instance, on dataset L5, the Transformer-based baseline fails to achieve downlink CSI prediction, and correspondingly, WiFo-M2 provides no enhancement. This indicates that the supplementary features cannot compensate for fundamental algorithmic limitations in the baseline model.
V-D Ablation Study
In this subsection, we investigate the influence of various specialized designs in both the ContraSoM pre-training strategy and WiFo-M2 model, including temporal feature extrapolation (denoted by TFE) and diffusion-based LiDAR feature augmentation (denoted by DA). To highlight the crucial role of the ContraSoM strategy, we additionally evaluate WiFo-M2-Img whose weights are replaced by ImageNet-pre-trained ones, denoted as WiFo-M2-Img w/ INet Weight.
Due to page limits, we present ablation results for several representative PHY actions, as shown in Table IV. First, removing the temporal feature extrapolation function in WiFo-M2 leads to a performance drop, confirming that extrapolating fine-grained future features effectively bridges the sparsity of sensory data. The NMSE results reported are averaged over the L1–L7 datasets. The ablation study results demonstrate the effectiveness of each proposed component. Second, replacing the ContraSoM pretrained weights with generic ImageNet‑pretrained weights significantly degrades performance (e.g., from 9.63-9.63 to 7.55-7.55 on LPCCNet), highlighting the advantage of the supplementary channel features over general‑purpose visual features. Third, removing the DA design from WiFo-M2-LiDAR causes a performance decline, validating that the proposed feature‑space diffusion process enhances LiDAR representation learning.
TABLE IV: Results of ablation experiments.
CE: FCDAMP CI: LPCCNet CP: WiFo
WiFo-M2-Img 3.38-3.38 9.63-9.63 3.38-3.38
WiFo-M2-Img w/o TFE 3.23-3.23 8.87-8.87 3.30-3.30
WiFo-M2-Img w/ INet Weight 3.22-3.22 7.55-7.55 2.96-2.96
CE: FCDAMP CI: LPCCNet CP: WiFo
WiFo-M2-LiDAR 3.36-3.36 9.13-9.13 3.45-3.45
WiFo-M2-LiDAR w/o DA 3.27-3.27 8.53-8.53 3.27-3.27
V-E Cross-Scenario Generalization
In this section, we evaluate the cross-scenario generalization ability of WiFo-M2 on the datasets C1, V1, and S1, which are not seen in the pre-training datasets.
V-E1 S1 Dataset
The S1 dataset provides measured image and LiDAR point cloud data, as well as the receive power with beams. Due to the absence of CSI, only the performance of BP action is evaluated on S1 dataset. Note that BP action on S1 dataset is relatively straightforward due to the small 16-element antenna array. The Vision-BP baseline achieves Top-1/3/5 accuracies of 91.0%, 98.7%, and 100%, while the MM-BP scheme reaches 92.2%, 100%, and 100%. Notably, the WiFo-M2-img model, with its pre-trained backbone frozen and only a final MLP layer fine-tuned, attains 91.0%, 100%, and 100%, matching or surpassing the strong baselines. This demonstrates that the representations learned by WiFo-M2 successfully generalize to the unseen S1 scenario.
V-E2 C1 Dataset
Table V summarizes the cross-scenario generalization performance of WiFo-M2 variants when applied to CE, CI, and CP actions on the unseen C1 dataset. As can be seen, the WiFo-M2 model with multi-modality provides consistent and often the best performance gains across most schemes, achieving the lowest NMSE for both CE schemes and one CP scheme (WiFo). Importantly, the WiFo-M2-Img consistently outperforms its ImageNet-pre-trained counterpart, confirming the effectiveness of the proposed ContraSoM strategy. BP results exhibit the same trend. Using only LiDAR, the MM‑BP-LiDAR baseline achieves top‑1/3/5 accuracy of $69.4\%$, $86.0\%$, and $86.0\%$, while WiFo-M2-LiDAR attains $73.1\%$, $90.3\%$, and $99.5\%$. With image alone, the vision-BP baseline delivers $78.5\%$, $88.2\%$, and $90.3\%$, against $77.4\%$, $85.5\%$, and $89.2\%$ for WiFo-M2-Img. Under multi-modal fusion, MM-BP scheme peaks at $88.7\%$, $99.5\%$, and $100\%$, whereas WiFo-M2 achieves $83.9\%$, $91.4\%$, and $93.0\%$.
TABLE V: Evaluation of the enhancement effects of WiFo-M2 models on various PHY actions on C1 dataset. Values in bold indicate the best performance.
Schemes PHY Actions CE: CENN CE: FCDAMP CI: LPCCNet CI: SR-CI CP: WiFo CP: Transformer
Baseline 4.74-4.74 5.30-5.30 16.03-16.03 8.68-8.68 10.74-10.74 14.06-14.06
Baseline + WiFo-M2-Img w/ INet Weight 4.87-4.87 5.38-5.38 16.73-16.73 7.29-7.29 11.25-11.25 13.58-13.58
Baseline + WiFo-M2-Img w/ ContraSoM Weight 4.96-4.96 5.43-5.43 17.21-17.21 9.19-9.19 11.58-11.58 14.23-14.23
Baseline + WiFo-M2-LiDAR 4.77-4.77 5.43-5.43 17.24-17.24 8.59-8.59 11.45-11.45 14.28-14.28
Baseline + WiFo-M2 4.98-4.98 5.49-5.49 17.29-17.29 9.11-9.11 12.21-12.21 14.26-14.26
TABLE VI: Evaluation of the enhancement effects of WiFo-M2-Img model on various PHY actions on V1 dataset. Values in bold indicate the best performance.
Schemes PHY Actions CE: CENN CE: FCDAMP CI: LPCCNet CI: SR-CI CP: WiFo CP: Transformer
Baseline 3.09-3.09 3.17-3.17 8.81-8.81 3.85-3.85 5.23-5.23 -
Baseline + WiFo-M2-Img w/ INet Weight 2.33-2.33 4.28-4.28 9.52-9.52 4.32-4.32 5.02-5.02 -
Baseline + WiFo-M2-Img w/ ContraSoM Weight 3.41-3.41 5.17-5.17 10.46-10.46 4.35-4.35 5.55-5.55 -
V-E3 V1 Dataset
The V1 dataset provides only images and CSI data. Hence, we focus on the performance of WiFo-M2-Img. In BP action, the WiFo-M2-Img-based scheme achieves top‑1/3/5 accuracy of $67.4\%$, $91.2\%$, and $98.1\%$, outperforming Vision‑BP ($65.5\%$, $82.4\%$, $92.3\%$) and closely matching the performance of the MM‑BP scheme ($71.3\%$, $87.4\%$, $99.2\%$) which uses only its image branch. Table VI presents the enhancement effect of WiFo-M2-Img on other baselines, demonstrating consistent performance improvements across all baselines, with the exception of the Transformer-based CP scheme, which failed to converge. In the same way, the weights derived from the ContraSoM strategy play a significant role.
TABLE VII: Additional fine-tunable parameters introduced by applying WiFo-M2 to different PHY actions.
PHY Actions CE: CENN CE: FCDAMP CI: LPCCNet BP
Parameters (M) 0.13 0.26 0.09 0.18 0.035 0.051 0.067 0.13
PHY Actions CI: SR-CI CP: WiFo CP: Transformer
Parameters (M) 0.017 0.034 0.26 0.26 0.026 0.055
V-F Computational and Deployment Efficiency
To assess the practical deployment overhead of WiFo-M2, we evaluate its model complexity and inference latency. All experiments are conducted on a server equipped with dual 14-core Intel Xeon E5-2680 v4 CPUs and an NVIDIA GeForce RTX 4090 GPU, processing inputs with a batch size of $1$ to simulate real-time operation. WiFo-M2-LiDAR comprises only 0.070.07M parameters with an inference latency of 3.313.31ms, while WiFo-M2-Img contains 9.629.62M parameters with an inference latency of 6.046.04ms. When multiple modalities are available, the total inference can be executed in (at most) the sum of their latencies, i.e., 9.359.35ms, when processed serially. Even with this upper bound, the system stays well within a real-time budget: at $40$ fps a new multi-modal sample arrives every 2525ms, leaving 15.715.7ms after inference to exploit the freshly produced features. During those 9.359.35ms, the features predicted from the previous sample can still be used, which only requires the model’s output sequence to be extended to cover the 9.359.35ms inference window.
As another practical metric, Table VII presents the additional fine-tunable parameters introduced when WiFo-M2 is integrated into PHY actions. The parameter overhead remains extremely low, reaching at most 0.260.26M. This is due to the effectiveness and superiority of the ContraSoM pre-training strategy, which allows the WiFo-M2 backbone to remain frozen during adaptation and only requires the lightweight adapter to be fine-tuned together with the existing schemes. This further highlights the practical deployability of WiFo-M2.
VI Conclusions
In this paper, we have realized a framework where intelligent embodied network entities utilize multi-modal environment sensing to comprehensively refine their PHY actions in a plug-and-play manner. Central to our approach is a powerful foundation model, WiFo-M2, which is capable of extracting fine-grained OOB channel-aware features from sparse historical sensing inputs. This is enabled by ContraSoM, a contrastive pre-training scheme that aligns environment sensing with wireless channels in a shared latent space. To enhance representation robustness, we further incorporate modality-specific data augmentation techniques that apply stochastic pixel-space transformations for images and leverage a diffusion-based feature generation framework for LiDAR point clouds. Extensive experiments on typical PHY actions show that our proposed framework boosts network performance from on multiple fronts and generalizes well across diverse scenarios.